Space Research & Planetary Sciences Division Web Site

Planetary Imaging Group Web Site

CaSSIS News

9 March 2016

Launch - 5 days

Check out our fun video (produced by Ruth Ziethe) about how we built CaSSIS. 

For the more serious minded, we have made a video interview about CaSSIS in English and German.

 

 

8 March 2016

Press release about CaSSIS and the launch....  and in English.

4 March 2016

CaSSIS and TGO are now inside the rocket fairing.

Pictures from ESA's web site.

 

... and the fairing closed.

 

02 March 2016

Launch - 12 days

The covers are off.

18 February 2016

The SVT-2B test in Baikonur is the last switch-on and test prior to launch for CaSSIS. We ran a sequence including a 30 exposure push-frame in 3 colours with 2 dark areas at 1 ms exposure time and 400 ms repetition frequency (roughly mimicking a typical image acquisition for the prime mission). The sequence executed perfectly. The image below simulates the whole detector and Antoine has put the framelets from an actual exposure into this frame so you can see where on the chip the individual framelets were read from. The large framelets are (bottom to top) PAN, RED, and NIR. 

10 February 2016

Electrical Interface Model (test bed for the spacecraft/instrument interface) updated with the latest version of the Flight Software.

8 February 2016

Commands delivered for the next System Verification Test in Baikonur.

1 February 2016

43 days to launch.

We are just completing the procedure for commissioning the instrument in flight. The first sequences will be run on 7 April in the current planning. Around 12 April, we will have the first pointing campaign with images of specified targets. We are currently planning a star (theta Crt) as our first target. A starfield (Area 104 of the Landolt catalogue) will be our second target for geometric calibration.

The exposure times have been calculated for both targets but with a little margin. All of our checks with integrating sphere data and with knowledge of the system gain suggest that our exposure times should be good to 10% or so.

The pointing requirements have been delivered to ESOC and the first versions of the command files will be sent to ESOC this evening.

 

1 February 2016

We are just setting the default window sizes and positions on the detector. We need to specify the rows that will be returned to give us the four filtered images. The filters are fixed above the detector. Antoine has looked at this and provided the numbers. They will be

354-633  PAN

712-967  RED

1048-1303  NIR

1409-1664  BLU

 

1 February 2016

TGO and CaSSIS in Baikonur. Below are a couple of pictures of the spacecraft and the red tag cover on CaSSIS. We have operated the instrument a couple of times to check out problems. Things look good however. We had a Spacewire communication issue which turned out to be a feature and not a bug! Mostly minor stuff.

Above is the instrument (nadir-pointing) panel.

Above is a close-up of the CaSSIS CRU (left) and the E-box (right). 

 

30 December 2015

CaSSIS switched on in Baikonur. Seems OK but we need to check in detail.

22 December 2015

TGO transported to Baikonur.

3 December 2015

System Verification Test (SVT) completed. CaSSIS commanded from Darmstadt by ESOC. Instrument in Cannes on the spacecraft. We acquired 44 images and 1 incomplete image. Couple of bugs were found but it looks pretty good. Example image is below. Yes. It is dark. The aperture cover was on!

 

25 November 2015

Press conference to show spacecraft is ready to ship to Baikonur.

..... and a quote or two ......

 

13 November 2015

 

CaSSIS on the spacecraft!!

 

No pictures ... not allowed ..... CaSSIS is on the spacecraft.

9 November 2015

The cans are in Cannes .... I feel a tune coming on ....

 

8 November 2015

Packed and in the truck ......

 

and..............................

CaSSIS on its way to Cannes

Press release and Video with clips of tests (not a nice video but good for TV!)

7 November 2015

Packing. Acquired with our monitoring camera while the PI sits at home, drinking tea.

 

..... and one last check of the dust contamination.

 

 

6 November 2015

ESA are just leaving after the Delivery Review Board meeting. The meeting closed with us being given permission to ship to Cannes!!!

This is not acceptance to fly. That has to go through the flight acceptance review and we still have to justify some waivers and ..... stuff. But the truck leaves on Monday morning as planned.  

6 November 2015

Metrology. Centre of gravity, mass and moment of inertia being measured.

 

Update: Mass is 18.009 kg (300 g over our original guess = 1.7%... not bad.)

6 November 2015

Those of you who have read all about CaSSIS will know that it is a push-frame camera. We take 2D pictures every 350-400 ms as the spacecraft flies over the planet's surface. Now that means timing is important. At 400 km above the surface in a circular orbit, that framing rate must be one image every 367 milliseconds. But when we command that, do we get what we asked for?

We devised a little test with an LED and an arbitrary waveform generator operating in pulse mode. We slightly offset the pulsing with respect to 367 ms so that we could see the images changing from bright to dark over a period of 40 images. And sure enough, it showed that the instrument is roughly correct. I say roughly because the error bar on the number is a couple of milliseconds. But we have allowed about 10 milliseconds in the requirements - it is the overlap that we need to stitch the individual framelets together. OK. It's complicated but ...... The test shows that CaSSIS has no timing issues.

4 November 2015

The CaSSIS team in Bern (minus 4 guys who were busy elsewhere and Harry without whom .......).

4 November 2015

The final major test was completed on Sunday. The modification to the rotation drive worked and the telescope now turns 180 degrees in 16 seconds at -27C.

A final focus check with an interferometer while CaSSIS was in the vacuum chamber suggests that the focus is where we wanted it within measurement error.  

Looks like we are good to go.

The intention is now to deliver the instrument to Thales-Alenia Space in Cannes on Monday (9 November). Before that can happen we still have a load of admin to complete but the instrument is good to go.

 

4 November 2015

Stupid things we have to do ....... number 341 .... Verifying that a protection cover can be lifted by a crane without the hook coming off. 

 

1 November 2015

09:00 on a Sunday morning. Getting ready to change the orientation of CaSSIS so that full functional tests can be carried out over the full temperature range during thermal cycling. (Image of the CHEOPS lab. with the calibration set-up in front of the chamber.

30 October 2015

Status

CaSSIS is currently in thermal balance. With the MLI everything is very, very slow. The instrument takes more than 36 hours to reach thermal equilibrium. The heaters seem to be switching correctly and we have completed a bunch of tests on the detector at operational temperature (0 C). With the exposure times we are using there will be almost no dark current at this temperature.

We have also performed some relative spectral response tests. These are not quite completed but a first look shows that in PAN and RED, the spectral response is very close to predict. We still need to look in the NIR and BLU channels.

Mounting of the MLI has also had an effect on our EMC problem. It is still possible that we will not be able to image during ELECTRA passes but we can be in standby and so we can image within a few minutes of an ELECTRA pass ending. CRISM and MCS on MRO do something similar.

One concern coming from the thermal balance test is that support structure (containing the mechanism) is colder than predicted. (The conductivity through the feet is lower than modelled.) So we have to be sure that the mechanism will rotate at -20 C.  

Future planning

ESA want us in Cannes before 17:00 on 9 November 2015. We need to complete thermal cycling before then and perform metrology verification (that's centre of gravity and moment of inertia measurements). During the thermal cycling, we will do the software acceptance.

For the software, we have been running version 0.14 provided by the Hungarian team. This version seems to be stable and has full functionality except for the compressor. 

 

 

29 October 2015

Marc Erismann checking the CRU transport bell.

26 October 2015

CaSSIS is now covered with multi-layer insulation and is in the vacuum chamber for thermal balance testing.

 

22 October 2015

Vibration successfully completed.

 

YES!!!!!!!!!!!!!!!!!!!!!!!

21 October 2015

09:30 Y-axis sine acceptance completed.

14:15 Y-axis random acceptance completed. System damping more than assumed. Looks OK so far.

19:00 X-axis completed.

20 October 2015

Update. 1st resonance search (Y axis) completed. First frequency at 107 Hz.

 

20 October 2015

How about another pretty picture of the system just before we bag it for the vibration test.

17 October 2015

We have had a few problems with the flight software recently. But now most of the required elements are in (including a preliminary compression algorithm) and the boot is stable. We can also patch the software reliably. There are still a bunch of functional tests to complete with the latest version including image speed. Once debugged, we will deliver CaSSIS with this version. 

16 October 2015

Instrument is back from EMC testing at Montena. One issue. Our SpaceWire line between the electronics box and the detector proximity electronics radiates quite strongly. Normally you wouldn't worry too much but the frequency is quite close to the frequency that the NASA provider Electra communication system uses for talking to the rovers on the Martian surface. We think we are OK because the frequency is sufficiently different but ESA is checking with JPL just to be sure.

On another topic, we now know that the gear mechanism didn't block at low temperature (see below)..... it just got too stiff for the little motor to turn the gear and we think we know why. We are just checking a solution for this before we go to the vibration facility.

 

11 October 2015

CaSSIS left the vacuum chamber after an initial test run. This was not entirely successful with the drive mechanism getting blocked at low temperature. This was partially caused by gravity. So we have taken the instrument out and will improve the set-up before performing the final TV qualification runs.

In the meantime, the instrument has had the stand-offs attached for the multi-layer insulation. You can just see the small brown pins sticking out from the structure in the picture below.

CaSSIS is now going to a company called Montena which performs tests on the instrument for electromagnetic cleanliness. It will be back here on Thursday night and go to the vibration facility before returning to the TV chamber.

29 September 2015

In the vacuum chamber with bake-out ready to start.

From the outside, the chamber looks like this .....

 

28 September 2015

We are trying to get to the vibration test but there are some issues because the tests of the spacecraft are indicating a slightly different vibration spectrum to the one we expected. Hence we are having to do some more calculations to prove that CaSSIS can handle it. We can produce some fun pictures from this. There are movies of the different modes which I will upload as soon as I figure out how to do it in typo3. :(

 

 

22 September 2015

CaSSIS lives!

The EGSE displaying the first image out of the CaSSIS system in flight configuration. All sub-systems have checked out successfully and can be commanded through the flight software.

Kudos to the electrical engineering team for completing a complex electrical integration in less than 24 hours. 

22 September 2015

The CaSSIS system in flight configuration (ok..... without MLI) for the first time.

... and bagged for transport down to B28 for full functional test and EMC testing tomorrow.

20 September 2015

To the sounds of "We Are The Champions", the telescope was mounted on the rotation drive last night. We are cabling and getting ready for the Full Functional Test.

 

18 September 2015

The integrated proximity electronics box and the twist capsule (the cable management system for making sure cables don't get snagged when we rotate) are mounted in the centre of the rotation bearing.  We start with positioning.

The PE is then mounted to the bipods. These bipods are important because during vibration there will be some amplification of the vibrational loads here. Something we have to watch for in the final test.

Completed. Now ready for wiring and mounting the telescope.

 

18 September 2015

Wiring up of the proximity electronics and the twist capsule.

15 September 2015

We have to do some very funny things sometimes..... like measuring the capacitance of an optical cube on the telescope....... no.... I have no idea either.

 

14 September 2015

Resonance search for the Camera Rotation Unit. Lead engineers Servonet and Gerber checking the data.

The CRU structure on the shaker with a dummy telescope.

Resonance search indicates a first eigenfrequency EXACTLY where Martin Rieder predicted it. No surprises. We even went to -3 dB for a short random shake with this. That's all fantastic. The slight downside of being right is that we also saw the predicted amplification for the proximity electronics (which sits in the central bearing of the rotation mechanism). We will have to be very careful with that. Now we complete the integration in preparation for the real shake test on Friday.

14 September 2015

We acquired some data on Friday night that was designed to test the linearity of the detector system. Vicky Roloff has done a quick analysis of that data and it looks surprisingly good.

Please note that the plots we show on the site are preliminary after a very rapid reduction step. All need to be checked before a final calibration report is produced.

11 September 2015

Pre-integration of the flight CRU before we look for resonances.

10 September 2015

Relative spectral response measurement acquired today showing the response of the instrument through the 4 filters when illuminated by monochromatic light at 5 nm resolution. We have a predict for this on this page. When comparing, note that below we have an absolute response - here it is relative in each filter. This was a very first plot to demonstrate we have the data. A proper calibration will follow.

9 September 2015

Start of the mechanism integration into the AlBeMet structure. Sascha Hausler and Matthias Brändli in the clean room.

One problem left to solve and that is connected to the drive shaft alignment with planetary gear. A small misalignment causes a loss of torque. We need to fix it. Current rotation rate to gain stereo is 15.6 seconds. Our requirement was 15.0 seconds. We have a bit margin so that's how it will fly!

9 September 2015

Focus - job done.

3 September 2015

Rotation end switch ready for integration.

2 September 2015

Rotation drive ready to have a pre-shake with the rotation ring.

2 September 2015

AM: ELU Y-axis shake complete. Limited functional test running but looks good so far.

Later: ELU shake completed. Functional test to be performed.

31 August 2015

ELU shake test preparations. One axis resonance search completed. 190 Hz first eigenfrequency.

 

31 August 2015

Preliminary flat-fields and linearity test completed. Here's a flat.

 

28 August 2015

One integrated ELU. The box is integrated and will go to the shaker probably on Tuesday.

28 August 2015

We joked below about a Hubble Space Telescope moment. And we had one and it is one reason we have been quiet for the past week. We tried to cross-check our mirror collimator with a lens collimator that we also have.... and got different results! So we have had to cycle everything right back to check what was going on. We have spent a lot of time looking at these .......

In the end, we have borrowed an interferometer from Thomas Feurer at the Institute of Applied Physics and that is showing that we are not quite in focus but close. So we are back on some sort of track but we will have to go back to re-shim the focal plane one more time.

P.S. If you want to build your own CaSSIS, don't try to get away with cheap calibration equipment. *makes mental note*

22 August 2015

Optical system moved with the lens collimator now pointing at the instrument ready for fine alignment before we start pumping.

Marc next to his work!

22 August 2015

The optical set-up is now being moved so that we can follow the effect on the focus of the drying out of the carbon-fibre structure in vacuo. The focus position should shift by about 46 microns. We have pre-compensated for this.

We are still a little nervous about the focus and so we will recheck our collimator with the Institute for Applied Physics's interferometer on Tuesday - just in case.

21 August 2015

The telescope with integrated focal plane have been moved to the vacuum chamber.

20 August 2015

The telescope has the detector in the focal plane. We will recheck this with a different collimator in a couple of days (to avoid a Hubble Space Telescope like issue ..... although Nicollier might fancy a flight to Mars). The system is now being transferred to a vacuum chamber in the calibration lab. for further tests.

We had one thing that didn't work so well. We tried a geometric distortion measurement which gave rather poor data. We would normally expect to get our best data using starfields in flight but we wanted something on ground just to demonstrate that we have a good model. Na ja. Next time.

In the meantime, the electronics guys have been busy getting the ground support equipment to display the telemetry from the flight software so that we have real-time data display when the electronics unit is hooked up to the telescope. 

They are also testing the hardware and software at high temperature to check for problems.

18 August 2015

The detector is torqued and in the correct rotation position. The radiator has been mounted.

Now checking the focus again at several positions to be sure we are OK.

18 August 2015

The focal plane is mounted and appears to be about 20 microns from the optimum position. During torqueing we got an 0.8 deg rotation of the focal plane - which is a little bit too much. So we will correct that this morning and mount the radiator before the final verification of the position of the FPA. 

15 August 2015

We are attaching a radiator to the focal plane assembly to cool the detector down. We would like to get rid of the dark current which is quite significant at room temperature. In space we will try to operate at 0 C. The integration to the FPA will be next week.

15 August 2015

The calculated focal length looks good and we can see in the residuals (not shown here) that the system has a slight distortion. We will be able to quantify this better later.

The UBE team would like to thank Vania Da Deppo from Padova for her support over the last 3 weeks. One of the best space optics specialists in the business. Here she is in the lab. with Lisa (who is standing behind the telescope).

 

15 August 2015

After a lot of checking of numbers and discussion, we will now produce the shims to position the detector at the optimum focus. Our target is -326 (+/- 50) microns taking into account the fact that the structure is carbon-fibre and will dry out in vacuum.

We decided to leave out the BLU filter data in optimizing the focus. We expect lower signal in the BLU (Mars is red in case you didn't know) and so we will probably bin the BLU data 2 x 2 in most cases. As the point spread function in BLU is in any case slightly inferior to the others (from the optical design and manufacture), we go for sharp data in the other filters.

14 August 2015

In 67 days CaSSIS has to be on the road to Torino to catch a plane to Kazahkstan.

13 August 2015

A little touch of improvisation. Our geometric distortion check will be after fixing the focal plane. But nervous PIs like to know if the instrument field of view is OK a bit earlier than that. So we improvised a grid pattern and took a picture. We used an LED torch for the illumination so that we would get good signal in the PAN and the BLU filter on the two ends of the detector (normal light bulbs give much more signal in the central RED and NIR filters). This worked well. You can see the grid (lined paper) right up to the edge of the filter mask. All this is giving us confidence that the optics and detector are doing what we expect.

13 August 2015

Some things are sent to try us. A little bit of kapton tape we were using to help us protect one of the mirrors in case we dropped a screw became unstuck and just cut a tiny fraction of the beam in the telescope. We caught it eventually and removed it. We checked it by taking a picture into the telescope. You can see here the filters in the focal plane very well indicating that we should be unvignetted. You can also see the field stop surrounding the detector which helps reduce straylight. You can also see the image of the photographer reflected off the filters. The squiggly stuff at the top are the heating elements used to stabilize the telescope temperature.

 

12 August 2015

We have to check whether electronics components get too hot when they are run in vacuum. You have no air to conduct heat away in space and so things can get hot really fast. Our thermal IR pictures show that the heat sinks we are using to keep the power supply voltage converters cool work pretty well.

10 August 2015

The electronics boards are coated and in their frames ready for final integration. However, first our Hungarian colleagues will arrive tomorrow for flight software debug. So we stay open for the moment just in case we have a hardware issue.

10 August 2015

Antoine has done a quick analysis of a dataset we got on Friday with the focal plane roughly in place. He used a 2D Gaussian to fit the point spread function as we did a through focus test. Although the focal plane is not yet fixed, it should show us what is possible with the system when everything is bolted down. You can see here that the system has astigmatism (the x and the y best foci are in different positions of the focal plane - as expected) and a nice round PSF is expected where the two curves cross. That's our target. The full width half maximum will then be <=1.3 px. Why <? Well, our pinhole is finite in size (10 micron) and so the PSF should be a little smaller in reality.

8 August 2015

USAF target imaged by the telescope. You can see the individual filter strips easily here. Top is the blue filter (almost dark because the source has a low temperature), NIR, red and panchromatic (yellow/red) at the bottom.

 

In the electronics lab. the power converter has been placed in its frame ready for integration into the complete electronics box. The Polish team are working with lead electrical engineer, Claudio Zimmermann, taking a picture.

7 August 2015

PI in the lab! The hardware got a prime visit today.

 

7 August 2015

Inspection of the electronics boards by ESA completed. 

They are now in coating and will be integrated into the electronics box this afternoon.

6 August 2015

The flight model power converter module has just arrived. Our Polish colleagues from SRC, Warsaw, are here to supervise the integration with the other flight boards. We are not expecting major problems because they already delivered us an engineering board which is very similar. However, it has to be tested in the next 2-3 days.

5 August 2015

The FPA is now attached to the telescope. The mechanical alignment and focussing procedure are starting now.

and here we have the first image from the flight model system (telescope and detector).

 

The little spot is the pinhole of our collimator. It shows that the instrument is close to but not quite in focus. We also confirmed a prediction about the misalignment between the optical axis of the telescope and the optical axis of the collimator. We can now move towards optimizing the system. Will probably take a couple of days to do that.

We have now just confirmed that the asymmetry in the PSF you see occurs when the detector is not in the focal plane. When we shift the position to get a focus, we get a nice round spot. Now we start getting the detector to the correct position before locking it in place.

4 August 2015

The integration team needed a break.

4 August 2015

PFM support structure delivered. This is a gold-coated piece of AlBeMet. It will form the support structure for the rotation mechanism. It now needs to be cleaned before we begin the integration of the bearing.

 

It weighs about a kilo.

4 August 2015

We did one last check of the Cable Management System (Twist Capsule) by hooking up the qualification model to the detector proximity electronics to verify the Space Wire connection. This worked fine and so we will finalize assembly of the Flight Model Twist Capsule. The proximity electronics box will now go to the integration room to drive the focal plane when it is connected to the telescope.

2 August 2015

Teams that work over the weekend require "support". 

1 August 2015

The interface plate has been attached to the tripod ready to mount the FPA to the telescope. Why do we need this? Well the telescope is carbon-fibre and the FPA is aluminium - so we have a mismatch in the coefficients of thermal expansion. This hardware helps us solve that problem. 

31 July 2015

Telescope on the tower ready to be aligned with the collimator. FPA is currently being integrated with its radiator and will go on the telescope for the first time on Monday morning.

30 July 2015

Here's a fun one. The focal plane assembly has a filter mask in front of it with 4 different coloured windows.

These windows are separated by black masks which are designed to prevent cross-talk - that is light bouncing around from one field to another. In this way we keep the light in each filter bandpass as pure as possible. But of course we have to test it. 

 

A laser was fired at a filter window which should be opaque to that laser wavelength. And indeed it is. The semi-circles you can in the image here come from the diffraction pattern of the laser. The main beam of the laser is almost completely masked out (if you look really closely you just see it - the filter is not perfect although pretty close). You can also see the black mask. There is almost NO cross-talk. The scheme works.

NB. The broad band of light in the middle of the frame comes from the light in the lab.

 

 

29 July 2015

The flight motor has been driven by the flight electronics. We can turn with enough torque to complete the 180 deg rotation in 17 seconds (10% longer than our target but well within the 30 seconds we could live with).

We are getting science data out of the FPA through the flight software onto the ground support equipment and translating that directly into a PDS V4 compatible format (print out of part of it below). One or two issues but it works. 

     FSW_HEADER:

       #att CompressionRatio: 0

       #att CompressionRatio_Length: 1

       #att CompressionRatio_Tag: 5

       #att Data_Length: 1048576

       #att Data_Tag: 160

       #att Error: 0

       #att Error_Length: 4

       #att Error_Tag: 9

       #att ExposureTimestamp: 000000000311d976

       #att ExposureTimestamp_Length: 8

       #att ExposureTimestamp_Tag: 11

       #att FRAME_HEADER: 250B

       #att Offset: 0

       #att Offset_Length: 4

       #att Offset_Tag: 7

       #att PEHK: 02a000020000001f20002a1f400000000000000000000000000000000081000000000000000000000003000000c20abe0a120a00000000df0d00000000000000

       #att PEHK_Length: 64

 

       #att PEHK_Tag: 80

28 July 2015

RUAG's measurement of the PSF in the centre of the field. For comparison, our pixel size is 10 micron. Right at the edge of the field in the blue section of the detector, the PSF is a bit ratty but elsewhere it looks a lot like this.

The carbon-fibre structure of the telescope shrinks in space as it loses moisture. RUAG have computed this for us and defined where we have to put the active area of the detector.

 

 

 

22 July 2015

Telescope delivery!!!!!

The box leaves RUAG's premises in Zurich.

After unpacking in the CaSSIS lab. (103) at Uni Bern. (The 3 Stooges are Ghose, Gambicorti and Gerber!!)

 

You might recognize the PI and the PM on the other side of the plastic curtain outside the clean area.

.... and a close-up of the telescope itself.

 

21 July 2015

Couple of new pictures

Firstly, the guys from SELEX arrived with the Focal Plane Assembly and Proximity Electronics. Here they are in the Bern electronics lab. helping us get the system going.

And here is the data from the FPA on the unit tester showing that all is well.

 

 

 

19 July 2015

FSW

The flight software is RUNNING. This was a real weird one. From the report....

"CaSSIS DPM hardware design causes a clock shift on the SDRAM clock line, causing double word memory access to fail in rare conditions at ambient temperature in the laboratory. The error occurred only on when writing two subsequent 32-bit words with bit 47 set and bit 15 cleared."

Wow!

"The GR712RC allows to adjust the SDCLK delay in the purpose register. It can be used to advance the clock until it is within the required timing specifications related to the other SDRAM signal lines."

Double wow!! 

The above is the (very) short version. Once fixed the FSW ran stably for 13 hours! Great job from the hardware and software teams in Bern and Budapest. Thanks also to Gaisler (Sweden) for their support during de-bug.

Telescope

Test review board is tomorrow afternoon. Delivery on Wednesday morning (22 July).

Rotation mechanism

We are fighting issues with the gearing. We have vastly improved the integration process but we still have issues with the coating being damaged during rotation. We are now lubricating the gears and the bearings with a space-qualified lubricant. There is a small worry about this stuff outgassing and getting onto the mirrors. But, we think this will not be sooooooo.... bad because the mirrors are quite well shielded from the mechanism. Getting this right is delaying us but it has to work well. We will know more later this week.

And.......

We will submit an abstract to AGU this week. 

10 July 2015

Mandatory inspection point for the telescope at RUAG.

Further to the remark from 1 July, the resonance search showed that the internal baffle had a higher frequency than predicted. No internal baffle change needed. We go to full random vibe tomorrow.

 

1 July 2015

Things are moving forward very rapidly but there are niggling little problems in several places. None of them would normally be big issues but they all cost the one commodity we don't have ---- time.

  • The rotation mechanism has been integrated and runs. Unfortunately it is a bit stiffer than we planned. We could reduce the pre-load but then the instrument's first resonance frequency will go down and ESA/Thales won't like that. So we are looking to improve the gearing on the motor drive.
  • The telescope is getting ready for vibration. We have an issue with the internal baffle which, when it is shaken, couples to a resonance frequency in the telescope and that is definitely not good. We might have to stiffen the internal baffle but it is not quite sure yet. The models aren't perfect and we might be OK. We find out when RUAG does a resonance search for the first time at the end of the week.
  • The focal plane assembly delivered 3 weeks ago had a small defect and needed to be exchanged. The new FPA is due here before the end of the week.
  • The oscillator which is required to work with the FPA for the read-out was found to be too slow. We dodged a bullet because getting space-qualified oscillators usually takes ages. However, we could "borrow" one from another project. It works!
  • A number of other corrections/improvements to the electronics hardware have been made. So far so good.
  • The flight PCM is running late because of delays in component soldering. Our Polish colleagues are "busting a gut" to deal with this.
  • The flight software is still not stable. It runs but there seem to be issues with the real-time operating system. The guys are working hard to debug this but it is costing us time.
  • A lot of telemetry conversion routines have been written and we can translate the HK data and the raw (uncompressed) frames from the focal plane sub-system. This seems fine so far but we need representative tests.
  • We have a widget-based tool for producing commands to the instrument ready.
  • The calibration facility is looking good. Our mirror collimator is aligned and ready. The lens collimator (which we use to verify performance when we go to the big TV chamber) has also been re-aligned by TRIOPTICS in preparation.

Whew!


11 June 2015

  • The focal plane subsystem (including the focal plane assembly and the proximity electronics) has arrived. It is running and is now being powered via the CaSSIS power control module. Both the FPA and the PCM will get replaced by the final flight elements in 3 weeks or so. We have been commanding manually through the ground support equipment and it looks good.
  • The PCM, the rotation control module and the digital processing module (the latter two boards being flight) are connected and talking to each other. The hardware seems to be good.
  • The flight software has been loaded but is still not stable. It is stable when I sit next to it apparently. The guys have asked ESA to provide a seat for me on the spacecraft next to the instrument.
  • The telescope mirrors are mounted and aligned. We seem to be OK apart from a slight (1 mm) physical shift of the field of view which we have to compensate for. Not quite sure how yet but …. It will be OK. The focus looks good (<10 micron FWHM in the centre of field was the preliminary number). Delivery expected on 15 July.
  • Last parts for the structure (the heaters) are due here on 13 July. We begin integration of the rotation structure and mechanism in 1st week of July.

The next 8-10 weeks are going to be ….. ahhhh….. interesting. 

 

Planetary Imaging Group Web Site

Space Research & Planetary Sciences Division Web Site